# EPIC-SOUNDS: A Large-Scale Dataset of Actions that Sound

### v2

@@ I Introduction
In summary, we introduce EPIC-SOUNDS, a large-scale dataset of daily-life sounds, derived from the audio of EPIC-KITCHENS-100. EPIC-SOUNDS contains 78,366 categorised sound events spanning over 44 categories, as well as 39,187 non-categorised sound events, totalling 117,553 sound events across 100 hours of footage collected in 700 videos from 45 home kitchens. The sound classes are based on descriptions from only listening to audio, thus suitable for problems in acoustics such as audio/sound recognition and sound event detection.
In this paper, we begin by introducing the related works to EPIC-SOUNDS (Section II). We then introduce EPIC-SOUNDS (Section III) followed by detailing our data collection pipeline (Section IV). We present an extensive analysis of the interplay between audio and visual modalities, as well as the complexity of collecting sounds for material based collisions. (Section V). We release two challenges on EPIC-SOUNDS: sound recognition, classifying a sound event given its start and end time, and sound detection, both localising and classifying all sound events within an untrimmed video. Additionally, we provide strong baseline results for these challenges, using both audio-only and audio-visual approaches (Section VI). Finally, we conclude our work, reflecting on its contributions as well as providing an overview of the already achieved impact of EPIC-SOUNDS since its initial release (Section VII).
@@ II Related Work
Fine-grained audio-visual datasets. The PACS dataset [25] focuses on understanding the physical common sense attributes of objects shown in the video, which is similar to our ‘material’ based annotation procedure. However, these attributes are distinguished by 13.4K question-answer pairs; displaying the video with and without audio, and then querying a variety of physical properties. SSW60 [19] consists of 31K images, 3.8K audio and 5.4K videos of 60 species of birds, proposed to facilitate works on fine-grained categorization using audio-visual fusion. Both datasets do not contain temporal annotations of sounds.
@@ III EPIC-SOUNDS: dataset statistics
EPIC-SOUNDS. The dataset consists of 78,366 categorised temporal annotations with an average length of 4.9s, distributed across 44 classes. We match the train / validation / test splits from EPIC-KITCHENS-100, giving the per-class proportion across splits in Fig. 2 (left).
@@ IV-A Data collection of labelled temporal segments
Instructions to the annotators. We worked with 20 annotators hired from an annotation company. We use a customised version of the VIA tool [26] to gather the annotations,
shown in Figure 4.
@@
Annotators are asked to listen to the audio and detect any distinctive audio event. They then are instructed to mark the start and end time of each distinctive sound they hear. Each segment is then given a semantic label which best describes the annotator’s perception of the action associated with the audio event. We impose no restriction on the vocabulary used, so the annotators may describe this however they wish. As a guide, we provide a list of sound labels that commonly occur in daily life, which the annotator may refer to, though they are not required to explicitly choose from this list. We term a segment-label pair as an ‘audio annotation’. A second annotator performs quality assessment of the audio annotations produced by the first annotator, particularly focusing on any missed audio events.
For each unique label description, the VIA tool creates a separate time-line, effectively grouping sequences of the same event. Note that sound events can overlap in time. If two segments with the same label are less than 0.3s apart, we instruct the annotators to merge the two segments as we deem them to belong to the same event. Additionally, annotators are asked to identify consistent background sounds (or noise) that occur throughout a large portion of the audio (e.g. radio, fan or washing machine). The annotators were asked to tag these as ‘background’. The procedure described thus far resulted in the annotation of 556 distinct sound descriptions.
@@
Instead, they select the actual materials involved when viewing the video. We only retain visually-verified collision sounds – i.e. materials correctly perceived from the audio only, then verified from the visual observation. We choose all collision material labels for which at least 40 examples are present. As a result, abstract labels related to collision (e.g. ‘clang/clatter’, ‘put objects on surface’) are clustered into 24 sound categories describing the materials involved, such as C: metal-only, or C: plastic-wood. We use the letter C to indicate these are collision-based classes.
@@ IV-B Post-processing Annotations
For the training set, we utilise the overlaps between audio segments and visual segments to select the samples for reviewing. We deem the use of the visual labels acceptable for error correction, as the annotation process is complete. Thus, utilising the visual labels for post-processing no longer compromises the issues stated in Figure 1. We review all audio classes for which there exists a mapping to visual classes in EPIC-KITCHENS-100. We identify two types of mapping, trivial; the audio class itself already exists as a visual class e.g. ‘scrub’, and relational; the audio class does not exist as a visual class itself but can be semantically mapped to one or more of the visual classes, such as the audio class ‘click’ relating to the verb ‘turn on / off’ or the noun ‘light switch’. We consider all annotations not labelled as the audio class of interest, but which overlap with action clips containing its visual mappings. We then manually assess each overlapping annotation, through listening to the audio and determining the label’s correctness. We run this error checking cycle multiple times to ensure all incorrectly classified instances are accounted for.
Figure 6: Customised LISA [27] annotation interfaces used for manually cleaning the validation and test set in EPIC-SOUNDS. (a) Interface for manual checking the trimmed event labels. (b) Interface for correcting disagreeing samples from the manual checking stage. (c) Interface for choosing between different annotations.
@@
Figure 7: Overlapping visual classes with the audio classes in EPIC-SOUNDS. For each audio-class, the top-10 most frequent overlapping visual classes are shown, dividing the bar into the proportion of overlapping instances. Note that the plot is log-scaled but the bar division is linearly scaled.
Figure 8: Bar charts visualising overlapping audio instances (top) and silent visual instances (bottom).
@@ V Audio-Visual Analysis
V-A Temporal Overlap of Visual and Auditory Events
@@
When comparing audio to video labels, we reflect on our motivation in Figure 1. For each audio class, we study the overlap with visual classes. Figure 7 visualises the top-10 most frequent overlapping visual classes, per audio class. The figure shows the diversity of overlapping classes. We see strong associations between the visual class ‘open cupboard’ and audio classes ‘open / close’ and ‘footstep’, noting that in many cases the participants are walking-up to cupboards they are about to open. Note that we only display the top-10 frequently overlapping visual actions here.
@@
Figure 8 (top) shows the proportion audio events with an overlapping visual event. We hypothesise that an overlap typically represents an in-view audio-visual event (i.e. an audio event overlaps with a visual event). Conversely, no overlap implies the auditory event corresponds to an out-of-view visual event. The figure thus showcases the proportion of in-view audio-visual events for each audio class. We see that the top three audio classes with out-of-view visual actions are: footstep (28.5%), click (21.1%) and plastic-marble collision (20.1%). In these instances, the classes relate to sounds produced by visual actions which happen off-screen, or are occasionally deemed trivial, such as placing an object or turning on the hob, resulting in missed visual annotations while still producing distinctive auditory signals. Conversely, the three classes which most frequently overlap with visual classes are zip (100%), sizzling / boiling (97.1%) and hoover / fan (97.0%). For zip, these relate to clear visual actions that are frequently annotated (opening/closing a bag, or cleaning a surface), whereas sizzling / boiling and hoover / fan relate to long-form audio which occurs during multiple visual activities, such as pan frying food when completing other steps of a recipe, or when an extractor fan is switched on while cooking.
@@
We also visualise the proportion of silent instances of visual actions in Figure 8 (bottom). These are actions that do not overlap with an audio event. Here we plot a representative sample of visual actions, across various proportions of silent instances. We see the top-3 video classes that have no overlap with audio (0-to-1) are: insert mixture (84.0%), fold dough (50.0%) and hang cloth (41.4%). Indeed these are actions that seldom produce a discernible sound. Conversely, we see that the top-3 visual actions with auditory counterparts are wash spoon (98.2%), wash glass (97.0%) and pour water (95.3%) where clearly the water will produce clear audio signals that are easily recognisable.
@@
We investigate repeated sounds across a single visual action. On average, we find the top-3 many-to-1 audio-to-video classes that contain repeated audio sounds (on average) are: cut / chop (2.28-to-1), beep (1.47-to-1), metal / wood collision (1.24-to-1), these relate to actions that have a ‘stop-start’ pattern e.g. pauses between chops, button presses on an appliance, or between repetitively moving items in a cutlery drawer or sink.
@@ V-C Material Analysis
We compute two metrics per material and report the results in Figure 10. Red bars represent the ratio of materials recognised purely from audio to materials verified from vision (Of those visually verified as material X, how many were pre-labelled as material X?). Blue bars represent the ratio of material recognised purely from audio then verified by vision (Of those labelled as material X only with audio, how many were visually verified as material X?). Results show that metal performs highly on both metrics due to its distinctive, resonant sound when struck, which is easily distinguishable from other materials. Cloth exhibits low performance, especially on the ratio of visually verified cloth sound to the cloth sound recognized only with audio. Cloth typically produces muffled, soft sounds upon impact, which are less characteristic and more easily confused with other materials or ambient noise.
@@ VI Challenges and Baseline Results
In this section, we define and experiment on two challenges: sound recognition and sound detection. For each challenge, we evaluate both audio-only models as well as audio-visual models to measure the complementary nature of the visual modality.
@@ VI-A Challenge Definitions
Sound Recognition. Given an audio segment $S^{i}=[t^{s}_{i},t^{e}_{i}]$, we aim to classify the ongoing sound event within the segment $c^{s}_{i}\in C$, where $C$ is the 44 classes in EPIC-SOUNDS. Note that for this challenge, the start and end time of all segments are known, meaning that all start and end times are given during inference, and the model is only expected to classify the sound. To assess this challenge, we report the top-1 and top-5 accuracy, as well as mean average precision (mAP), mean area under ROC curve (mAUC), and mean per class accuracy (mCA) during inference.
@@ VI-B Audio-only Sound Recognition
Baselines. We train and evaluate the Auditory SlowFast (ASF) [29] and Self-Supervised Audio Spectrogram Transformer (SSAST) [28] audio encoder networks, with both a linear probe, i.e. by freezing the model weights and only training the last classification layer, and by fine-tuning. We also compare to a chance baseline. ASF is pretrained on VGG-Sound, and SSAST is pretrained on AudioSet and LibriSpeech [31].
Audio Processing. We follow the audio processing of [29] for extracting the input spectrograms for both models, noting that this outperformed the default audio processing of SSAST ($200\times 128$ spectrograms for 2s of audio, or $400\times 128$ for 4s of audio sampled at 16kHz). Namely, audio is resampled at 24kHz for both models. We randomly sample 2s of audio to create log-mel-spectrograms with 128 Mel bands. If the audio annotation is shorter than 2s we pad the produced spectrogram with its last column. We use a window and hop size of 10ms and 5ms respectively, resulting in a spectrogram of size $400\times 128$.
@@ VI-C Audio-Visual Sound Recognition
Results. We report the audio-visual results in Table III. Here, we see that the visual modality assists in audio-based interaction recognition, where both audio-visual baselines consistently improve across metrics. TIM outperforms MTCN on Top-1 accuracy on the validation set and mAP on the entire test set, with MTCN performing better in the remaining metrics. As MTCN is given neighbouring action start-end times during inference, it can exploit the relationships between neighbouring sound actions, specifically when considering repetitive sounds.
@@ VI-D Audio-only Sound Detection
Results. We report detection results in Table IV. In audio-only, TriDet outperforms ActionFormer by 0.1 average mAP on both the detection test set and the entire test set, whereas for the validation set, ActionFormer outperforms TriDet by 0.2 average mAP. Typically, ActionFormer appears to be more accurate with its regressed proposals, showing higher average precision at the strictest 0.5 IoU threshold for all splits. We show qualitative results for both TriDet and ActionFormer for two videos in Figure 12, along with a zoomed-in crop of a dense 20-second window in the video. Here, we show the predictions for the top-10 most frequent audio classes across the two videos, highlighting how the detection baselines can distinguish between multiple, potentially overlapping sound interactions.
@@
TABLE IV: Results of the Baseline Models on the EPIC-SOUNDS validation, detection test and entire test splits. We report the average precision at IoU thresholds $[0.1,0.2,0.3,0.4,0.5]$ and their average across all thresholds.

### v3

@@ I Introduction
In summary, we introduce EPIC-SOUNDS, a large-scale dataset of daily-life sounds, derived from the audio of EPIC-KITCHENS-100. EPIC-SOUNDS contains 78,366 categorised sound events spanning over 44 categories, as well as 39,187 non-categorised sound events, totalling 117,553 sound events across 100 hours of footage collected in 700 videos from 45 home kitchens. The sound classes are based on descriptions from only listening to audio, thus suitable for problems in acoustics such as sound recognition and sound event detection.
In this paper, we begin by introducing the related works to EPIC-SOUNDS (Section II). We then introduce EPIC-SOUNDS (Section III) followed by detailing our data collection pipeline (Section IV). We present an extensive analysis of the interplay between audio and visual modalities, as well as the complexity of collecting sounds for material based collisions. (Section V). We release two challenges on EPIC-SOUNDS: sound recognition, classifying a sound event given its start and end time, and sound detection, both localising and classifying all sound events within an untrimmed video. Additionally, we provide strong baseline results for these challenges, using both audio-only and audio-visual approaches (Section VI). We conclude our work by reflecting on its contributions and provide an overview of the already achieved impact of EPIC-SOUNDS since it’s release (Section VII).
@@ II Related Work
Fine-grained audio-visual datasets. The PACS dataset [25] focuses on understanding the physical common sense attributes of objects shown in the video, similar to our ‘material’ based annotation procedure. However, these attributes are distinguished by 13.4K question-answer pairs; displaying the video with and without audio, and then querying a variety of physical properties. SSW60 [19] consists of 31K images, 3.8K audio and 5.4K videos of 60 species of birds, proposed to facilitate works on fine-grained categorisation using audio-visual fusion. Both datasets do not contain temporal annotations of sounds.
@@ III EPIC-SOUNDS: dataset statistics
EPIC-SOUNDS. The dataset consists of 78,366 categorised temporal annotations with an average length of 4.9s, distributed across 44 classes. We match the train / validation / test splits from EPIC-KITCHENS-100, giving the per-class proportion across splits in Figure 2 (left).
@@ IV-A Data collection of labelled temporal segments
Annotation process. We worked with 20 annotators hired from an annotation company. For each video segment, we use one annotator to give an initial set of audio annotators and a second annotator to check these. The first annotator listens to the audio and identifies distinctive audio events, marking their start and end times and assigning a semantic label that reflects the annotator’s interpretation of the action or source associated with the sound. Annotators use free-form vocabulary, though we provide a reference list of commonly occurring everyday sound labels to guide them which they may optionally select from. The second annotator performs quality control, e.g. reviewing the annotations for any missed events, and makes necessary corrections.
We use a customised version of the VIA tool [26] to gather the annotations. , shown in Figure 4. For each unique label description, the VIA tool creates a separate time-line, effectively grouping sequences of the same event. Note that sound events can overlap in time. If two segments with the same label are less than 0.3s apart, we instruct the annotators to merge the two segments as we deem them to belong to the same event. Additionally, annotators are asked to identify consistent background sounds (or noise) that occur throughout a large portion of the audio (e.g. radio, fan or washing machine). The annotators were asked to tag these as ‘background’. The procedure described thus far resulted in the annotation of 556 distinct sound descriptions.
@@
Instead, they select the actual materials involved when viewing the video. We only retain visually-verified collision sounds – i.e. materials correctly perceived from the audio only, then verified from the visual observation. We choose all collision material labels for which at least 40 examples are present. As a result, abstract labels related to collision (e.g. ‘clang/clatter’, ‘put objects on surface’) are clustered into 24 sound categories describing the materials involved, such as C: metal-only, or C: plastic-wood, where C indicates collision-based classes.
@@ IV-B Post-processing Annotations
For the training set, we utilise the overlaps between audio segments and visual segments to select the samples for reviewing. We deem the use of the visual labels acceptable for error correction, as the annotation process is complete. Thus, utilising the visual labels for post-processing no longer compromises the issues stated in Figure 1. We review all audio classes for which there exists a mapping to visual classes in EPIC-KITCHENS-100. We identify two types of mapping, trivial; the audio class itself already exists as a visual class e.g. ‘scrub’, and relational; the audio class does not exist as a visual class itself but can be semantically mapped to one or more of the visual classes, such as the audio class ‘click’ relating to the verb ‘turn on / off’ or the noun ‘light switch’.
We manually review all cases where an audio annotation overlaps with a different visual action – e.g. a ‘scrub’ audio class with one overlapping visual annotation of the class ‘open / close’. We use this additional filtering to correct the audio classes.
We run this error-checking cycle multiple times to ensure that all incorrectly classified instances are accounted for.
Figure 6: Customised LISA [27] annotation interfaces used for: (a) manually checking the trimmed event labels, (b) correcting disagreeing samples from the manual checking stage and (c) choosing between different annotations.
@@
Figure 7: Bar chart showing the diversity of temporal overlaps between visual and audio classes. The proportion of overlapping instances for each of the top-10 most frequent temporally overlapping visual classes divides the bar of each audio-class. Note, the plot is log-scaled, but bar division is linearly scaled.
Figure 8: Bar charts showing the proportion of audio classes that have no temporal overlap with visual classes i.e. no clear visual signal (top), and visual classes that have no temporal overlap with audio classes i.e. no audible signal and are thus considered silent visual actions (bottom).
@@ V Audio-Visual Analysis
V-A Overlap of Visual and Auditory Events
@@
When comparing audio to video labels, we reflect on our motivation in Figure 1. For each audio class, we study the temporal overlap with visual classes and make deductions based on such overlaps. Figure 7 visualises for each audio class the number of instances for the top-10 most frequent temporally overlapping visual classes. Individual bars are divided by the proportion of instances for each overlapping visual class. The figure shows the diversity of overlaps across the audio classes and implies that there are strong associations between visual and audio classes. For example, the visual class ‘open cupboard’ has frequent temporal overlap with audio classes ‘open / close’ or ‘footstep’, noting that, in many cases, participants walk towards cupboards they are about to open.
@@
Figure 8 (top) shows the proportion of audio events that have no temporal overlap with any visual event. We hypothesise that no temporal overlap implies that the auditory event corresponds to an out-of-view, or trivial visual event. Conversely, a temporal overlap between audio and visual classes typically represents an in-view audio-visual event. We see that the top three audio classes with the highest proportion of instances containing no temporal overlap with visual classes are: footstep (28.5%), click (21.1%) and plastic-marble collision (20.1%). In these instances, the classes relate to sounds produced by visual actions that typically happen off-screen, or are occasionally deemed trivial, such as placing an object or turning on the hob, resulting in missed visual annotations while still producing distinctive auditory signals. On the other hand, the three classes which most frequently have temporal overlap with visual classes are zip (100%), sizzling / boiling (97.1%) and hoover / fan (97.0%). For zip, these relate to clear visual actions that are frequently annotated (opening / closing a bag), whereas sizzling / boiling and hoover / fan relate to long-form audio which occurs during multiple visual activities, such as pan frying food whilst completing other steps of a recipe, or when an extractor fan is switched on while cooking.
@@
We also visualise the proportion of visual actions with no temporally overlapping audio event in Figure 8 (bottom). We hypothesise that no temporal overlap with audio classes corresponds to silent visual actions, whilst the presence of a temporal overlap indicates an audible visual action. Here we plot a representative sample of visual actions, across various proportions of silent instances. We see the top-3 visual classes with the highest proportion of no temporal overlap with audio classes (0-to-1) are: insert mixture (84.0%), fold dough (50.0%) and hang cloth (41.4%). Indeed these are actions that seldom produce a discernible sound. Conversely, we see that the top-3 visual actions with frequent temporal overlap with audio classes are wash spoon (98.2%), wash glass (97.0%) and pour water (95.3%) where clearly the water will produce clear audio signals that are easily recognisable.
@@
We investigate repeated sounds across a single visual action. We find the top-3 many-to-1 audio-to-video classes containing repeated audio sounds (on average) are: cut / chop (2.28-to-1), beep (1.47-to-1), metal / wood collision (1.24-to-1), these relate to actions that have a ‘stop-start’ pattern e.g. pauses between chops, button presses on an appliance, or between repetitively moving items in a cutlery drawer or sink.
@@ V-C Material Analysis
We compute two metrics per material and report the results in Figure 10. Red bars represent the ratio of materials recognised purely from audio to materials verified from vision (Of those visually verified as material X, how many were pre-labelled as material X?). Blue bars represent the ratio of material recognised purely from audio then verified by vision (Of those labelled as material X only with audio, how many were visually verified as material X?). Results show that metal performs highly on both metrics due to its distinctive, resonant sound when struck, which is easily distinguishable from other materials. Cloth exhibits low performance, especially on the ratio of visually verified cloth sound to the cloth sound recognised only with audio. Cloth typically produces muffled, soft sounds upon impact, which are less characteristic and more easily confused with other materials or ambient noise.
@@
Figure 11 shows the confusion matrix for material recognition from human annotators. Since annotators are capable of labelling multiple materials, we adopt the Multi-Label Confusion Matrix (MLCM) proposed by [28]. NPL (No Predicted Label) denotes cases where a material present in the ground truth was not predicted by the annotator. For instance, if the true label is metal / plastic and the annotator only predicts metal, the NPL column for plastic is incremented. Annotators most accurately identify metal (80%), paper (50%), and plastic (49%). In contrast, stone/marble (18%) and cloth (8%) are more difficult to recognise. Notably, glass is frequently misclassified as metal (42%) and other hard materials such as ceramic and plastic are often confused with metal as well.
@@
Figure 11: Material recognition confusion matrix. We adopted Multi-label confusion matrix proposed by [28]. NPL (No Predicted Label) indicates instances where the model failed to predict a material label that is present in the ground truth.
@@ VI Challenges and Baseline Results
In this section, we define and experiment on two challenges: sound recognition and sound detection. For each challenge, we evaluate both audio-only models and audio-visual models to measure the complementary nature of the visual modality.
@@ VI-A Challenge Definitions
Sound Recognition. Given an audio segment $S^{i}=[t^{s}_{i},t^{e}_{i}]$, we aim to classify the ongoing sound event within the segment $c^{s}_{i}\in C$, where $C$ is the 44 classes in EPIC-SOUNDS. Note that for this challenge, the start and end time of all segments are known, meaning that all start and end times are given during inference, and the model is only expected to classify the sound. To assess this challenge, we report: top-1 and top-5 accuracy, mean average precision (mAP), mean area under ROC curve (mAUC) and mean per class accuracy (mCA).
@@ VI-B Audio-only Sound Recognition
Baselines. We train and evaluate the Auditory SlowFast (ASF) [30] and Self-Supervised Audio Spectrogram Transformer (SSAST) [29] audio encoder networks, with both a linear probe, i.e. by freezing the model weights and only training the classification layer, and by fine-tuning. We also compare to a chance baseline. ASF is pretrained on VGG-Sound, and SSAST is pretrained on AudioSet and LibriSpeech [32].
Audio Processing. We follow the audio processing of [30] for extracting the input spectrograms for both models, noting that this outperformed the default processing of SSAST ($200\times 128$ spectrograms for 2s of audio, or $400\times 128$ for 4s of audio sampled at 16kHz). Namely, audio is resampled at 24kHz for both models. We randomly sample 2s of audio to create log-mel-spectrograms with 128 Mel bands. If the audio annotation is shorter than 2s we pad the produced spectrogram with its last column. We use a window and hop size of 10ms and 5ms respectively, resulting in a spectrogram of size $400\times 128$.
@@
We also perform a small-scale evaluation of human performance on audio-only sound recognition to compare human perception with computational models. We sample up to 10 clips per class from the EPIC-SOUNDS validation set, resulting in 412 clips nearly-balanced subset. Using the LISA interface (Figure 6), annotators (i.e. the paper’s authors) classify each clip using the EPIC-SOUNDS label set without access to video. Table IV shows the results. Humans achieve 20.8% accuracy, close to ASF (21.1%) and above SSAST (19.7%), aligning with the mCA in Table III due to balanced sampling. Notably, humans are especially effective at recognising non-collision sounds, achieving 36.2% accuracy, which surpasses ASF by 1.7% and SSAST by 3.5%.
TABLE IV: Human performance comparing to audio-only recognition baselines on a subset of the EPIC-SOUNDS validation set. We report the overall performance as well as performance on both collision and non-collision sounds.
Total Collision Non-collision
Human 20.8% 9.4% 36.2%
ASF [30] 21.1% 11.6% 34.5%
SSAST [29] 19.7% 9.8% 32.7%
@@ VI-C Audio-Visual Sound Recognition
Results. We report the audio-visual results in Table III. Here, we see that the visual modality assists in audio-based interaction recognition, where both audio-visual baselines consistently improve across metrics. TIM outperforms MTCN on Top-1 accuracy on the validation set and mAP on the entire test set, with MTCN performing better in the remaining metrics. As MTCN is given neighbouring action start-end times during inference, it can exploit the relationships between neighbouring sound actions, especially for repetitive sounds.
@@ VI-D Audio-only Sound Detection
Results. We report detection results in Table V. In audio-only, TriDet outperforms ActionFormer by 0.1 average mAP on both the detection test set and the entire test set, whereas for the validation set, ActionFormer outperforms TriDet by 0.2 average mAP. Typically, ActionFormer appears to be more accurate with its regressed proposals, showing higher average precision at the strictest 0.5 IoU threshold for all splits. We show qualitative results for both TriDet and ActionFormer for two videos in Figure 13, along with a zoomed-in crop of a dense 20-second window in the video. Here, we show the predictions for the top-10 most frequent audio classes across the two videos, highlighting how the detection baselines can distinguish between multiple, potentially overlapping sounds.
TABLE V: Results of the Baseline Models on the EPIC-SOUNDS validation, detection test and entire test splits. We report the average precision at IoU thresholds $[0.1,0.2,0.3,0.4,0.5]$ and their average across all thresholds.